OpenStudy (anonymous):

Which operation would isolate the variable in this equation. -74w=222 Choose one Divide both sides by 74 Divide both sides buy -74 Multiply both sides by -74 Add 74 to both sides.

OpenStudy (ash2326):

@PaigeMya What should we do to make \(\huge -74 w\) as \(\huge w\ ?\)

OpenStudy (anonymous):

I dont know that why i need help hahaha

OpenStudy (mertsj):

What does -74w mean?

OpenStudy (ash2326):

If we had 2w then what would you do to make it w? \[\huge 2w \to w\] ?

OpenStudy (anonymous):

Combine like terms?

OpenStudy (mertsj):

What does -74 w mean?

OpenStudy (ash2326):

Think like this. I have 4 and I want it to become 2 without using subtraction? What should I do?

OpenStudy (anonymous):

Uh subtract? 4-2?

OpenStudy (ash2326):

No I can't use subtraction. What should I do?

OpenStudy (anonymous):

I dont know haha sorry

OpenStudy (mertsj):

But she doesn't know she can't use subtraction. That is the point of the question.

OpenStudy (anonymous):

Oh i didnt know that..

OpenStudy (ash2326):

If we have 4 and I want 2. I'll use division \[\huge \frac{4}{2} \to 2\] Or If we had 2w I want w. I'll use division . Will divide by 2 \[\huge \frac{2w}{2} \to w\] Can you do it now???

OpenStudy (anonymous):

so 4/2 is 2 and then 2w/2 its 2w?

OpenStudy (ash2326):

\[\huge \frac{2w}{2}=\frac{\cancel 2w}{\cancel 2}=w\]

OpenStudy (anonymous):

Im really stupid guys hahaha ive missed a month of school

OpenStudy (ash2326):

No worries, Can you do the question now?

OpenStudy (anonymous):

I dont really know... i havent

OpenStudy (ash2326):

Try it. how will you get w from -73w

OpenStudy (anonymous):

I havent learned any of it yet. uh subtracting it from both sides

OpenStudy (anonymous):

or from one side but its actually -74w

OpenStudy (ash2326):

We have \[-74w=222\] to get \(\huge w\) we need to divide, can you try now?

OpenStudy (anonymous):

If i knew what to do next i wouldve ddone it but i dont s

OpenStudy (anonymous):

Someone please help!

OpenStudy (ash2326):

@PaigeMya We have \[\large -74 w=222\] to get w , divide both sides by -74 \[\large w=\frac{222}{-74}\] We get \[\large w=-3\]

OpenStudy (anonymous):

divide both by -74 is correct